The demand and supply for soft drinks are




given by Q = 10-P and Q = 2+2P, respectively. Suppose now the government imposes a per unit subsidy of $1.5 on the sellers.




a)what is the price received by the sellers after the subsidy has been imposed?

1
Expert's answer
2022-01-11T11:43:17-0500

We first rewrite the demand and supply equations as, "P=10-Q" and "P={Q-2\over2}" respectively. With "\$1.5" tax on sellers, the supply curve after tax is "P={Q-2\over2}+{1.5}={Q+1\over2}".

The new equilibrium quantity can be found by equating "P={Q+1\over2}"and "P=10-Q". So, "{Q+1\over2}=10-Q". This gives the new equilibrium quantity as, "Q_T={19\over3}".

The price sellers receive can now be obtained from the pre-tax supply equation as,

"P={Q_T-2\over2}={{19\over3}-2\over2}=2.17"

Therefore, the price received by the sellers after the subsidy has been imposed is 2.17.